Linoleic acid chemical formula C18H32O2 how many grams of linoleic are there in (175)0 of linoleic acid?


1
Expert's answer
2021-03-12T05:59:30-0500

M(C18H32O2) = 12"\times"18 + 1"\times"32 + 16"\times"2 = 280 (g / mol)

17.5 mol of linoleic acid contain 17.5"\times"280 = 4900 g of linoleic acid

Answer: m(C18H32O2) = 4900 g.